# Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?

### v1

@@ 2.2.4 Bicoherence calculation
The (auto-, also known as self-) bicoherence mendel1991tutorial ; he2022rossby is used to estimate the degree of phase-locking between 3 frequencies $f_{1}$, $f_{2}$, and $f_{3}=f_{1}+f_{2}$ in a given timeseries signal. Here, we use the classical definition of the norm-2 bicoherence:
$b^{2}(f_{1},f_{2})=\frac{|\sum_{n}F_{n}(f_{1})F_{n}(f_{2})F_{n}^{*}(f_{1}+f_{2})|^{2}}{\sum_{n}|F_{n}(f_{1})F_{n}(f_{2})|^{2}~\sum_{n}|F_{n}(f_{1}+f_{2})|^{2}},$ (5)
@@ 3.1 Tempelfjorden 2015 (T15)
In order to further investigate whether the high-frequency spectral peak and bicoherence peak can correspond to a nonlinear wave interaction caused by wave triads, we compute the real-world wave dispersion relation by applying the cross-spectral density analysis between the VN1005 and VN1007 IMU sensors. These are the 2 IMUs that are located furthest away from each other, which gives a higher relative accuracy in the phase-shift measurements. Due to the separation distance, phase aliasing is observed. This phase aliasing is removed by enforcing a smooth monotonic increase in $\Delta\phi$ for increasing frequencies, by applying suitable corrections by $2\pi$ when aliasing jumps take place. The results are presented in Fig. 3, with the dispersion relation from Eqn. (3) included for different effective sea ice thickness values and otherwise standard landfast ice mechanical characteristics (we use similar values as rabault2016measurements , i.e. $g=9.81$ m/s2, $E=2\times 10^{9}$ Pa, $\nu=0.3$, $\rho_{w}=1025$ kg/m3, $\rho_{i}=920$ kg/m3). As visible in Fig. 3, and in good agreement with the findings previously reported in sutherland2016observations , a significant deviation from the open water dispersion relation, which matches well with a thin-plate elastic dispersion relation with the same typical thickness as measured in the field, is observed early in the data (up to around segment 4). We show the data points obtained about the low signal-to-noise ratio higher frequency components by using the low CSC threshold as thin crosses, and we expect that these are particularly noisy and should, therefore, only be taken as an indication. Later on, the dispersion relation collapses back to the deep-water dispersion relation, which we interpret to correspond to the development of cracks through the ice cover. Cracks were indeed observed when the instruments were recovered, while they were not present when the instruments were deployed, consistent with this interpretation.
@@
Figure 3: Empirical dispersion relation obtained from a cross-spectral analysis between the IMUs VN1005 and VN1007. Clear deviation from the deep water dispersion relation, in good agreement with the thin elastic plate dispersion relation and with previous results sutherland2016observations , is observed until around segment 5. For later segments, the empirical dispersion relation matches the deep water one. Here, we show both the empirical dispersion relation obtained for the main PSD peak (large dots), and for the higher frequency components when these are present in the PSD (crosses; due to the low signal-to-noise ratio in the higher frequencies, these are obtained using a low value of the CSC threshold, and are therefore particularly noisy and should be taken only as an indication). The best-matching dispersion relation, which corresponds to an ice thickness of around 0.6 meters with the default ice parameters listed in the text (in good agreement with the field measurements reported in sutherland2016observations ), provides the $B_{\mathrm{ref}}$ value for the $B$ coefficient in Eqn. (2).
@@
Knowing that triads can be expected, one can also observe, back on Fig. 1, that there are two significant dips in the main peak of the PSD of the IMUs furthest in (VN1002 and VN1007) relatively to the PSD of the IMU furthest out (VN1005). These dips are observed consistently over several FFT points, for both IMUs further in the ice, at frequencies around 0.150.150.15Hz and 0.190.190.19Hz, which sum up to the typical frequency of the high-frequency peak (around 0.340.340.34Hz). These may be a signature of the energy that is being "stolen" from the main peak and transferred into the high-frequency peak by the triad interactions.
Figure 4: Wave triad diagram corresponding to the thin plate dispersion relation that best matches the empirical dispersion relation observed for early segments in Fig. 3 (i.e. elastic plate with thickness 0.6m). This dispersion relation can sustain the wave triad conditions in the frequency range of the high-frequency peak that we observe. For the specific value of $\mathbb{k}_{3}$ corresponding to the PSD and bicoherence high-frequency peak, sets of locii for $\mathbb{k}_{1}$ and $\mathbb{k}_{2}$ that match the main peak in the PSD are identified (red). In this and similar figures, $kx$ and $ky$ indicate the $x$ and $y$ components of a wave vector $\mathbb{k}=kx\mathbb{e}_{x}+ky\mathbb{e}_{y}$.
Since these data are obtained from field measurements, there are significant uncertainties on all parameters related to the ice properties (which are known to vary widely in time and space). In addition, the bicoherence peak is a relatively large "blob", and the PSD high-frequency peak is also relatively wide. Therefore, there is no single clear value for these parameters. Moreover, the "peak bicoherence triad" analyzed above corresponds only to the maximum of the bicoherence "blob". In reality, the spectra observed are broad-banded, and the high-frequency peak may rather be the consequence of a continuum of triad interactions than of a single triad. In particular, a full description of the dynamics would require integrating over all possible triads, akin to Hasselmann’s equation hasselmann1962non . However, this is beyond the scope of the present work, and to offer a convincing comparison between these more advanced descriptions and observations, one may need significantly more detailed field data including measurements from a large array of instruments.
@@ 3.3 Spectra over Iridium
In addition to these observations, we also considered the PSD spectra from deployments in the MIZ in the case where small broken ice floes were encountered. These include in particular the outermost buoys in the dataset 2021-03 of rabault2023dataset , see more descriptions in nilsen2021pc , and the outermost buoys in the datasets of muller2025distributed ; muller2025distributeddata . There, we could not find clear high-frequency peaks similar to those reported in Fig. 9. This further hints at the importance of having an ice sheet or a large enough ice floe to support significant flexural effects for observing the high-frequency peaks we report.
@@ 4 Discussion
In all cases, clear high-frequency peaks can be observed in some of the data segments. Since these high-frequency peaks are clearly above the noise background, are sometimes present and sometimes not (so they are not a systematic noise in the sensors and / or processing algorithms), and these can be observed over a range of instruments (VN100-based loggers, OMB-v2018 with a VN100 IMU, OMB-v2021 with an ISM330DHCX motion sensor, Sofar Spotter using GPS Doppler data) and deployments in various sea ice conditions and locations covering both the Arctic and the Antarctic, we can be confident that these are not noise nor a one-off coincidence, but a real recurring physical signal.
@@
A critical question to elucidate regarding the high-frequency spectral peaks is whether these correspond to bound or free waves. Bound waves, which canonical example is the higher-order components of the Stokes solution lamb1924hydrodynamics , are perturbations to the first-order wave solution that are present as the higher-order consequence of nonlinear terms. These can arise either from the effect of nonlinear terms on a single wave mode (as is the case with the higher order Stokes wave components), or as the effect of nonlinear terms on a superposition of wave modes (as the presence of nonlinear terms implies that the linear superposition of two or more wave modes results in an additional nonlinear residual term). Bound waves do not follow the dispersion relation, and are caused by and travel together with the main propagating wave, materializing as "disturbances" of the main wave. Free waves, on the contrary, propagate freely and follow the dispersion relation. Based on Fig. 3, it seems most likely that the high-frequency peak observed in the T15 dataset corresponds to a free wave that follows the dispersion relation. However, we acknowledge that the signal-to-noise ratio used there is low, so offering a definite confirmation that our high-frequency spectral peaks correspond to free waves would require additional observation and confirmation.
@@
While our observations seem to be compatible with the existence on nonlinear wave interaction and energy transfers between frequency modes through a wave triad mechanism, other mechanisms could also play a role and / or explain our observations. For example, zhao2017gap suggests that arrays of adjacent floating bodies (in our case, ice floes) can lead to nonlinear resonance and create high-frequency peaks in PSDs. However, this seems unlikely to be the cause of the observations in T15. Indeed, in this case, the high-frequency peak is observed when a flexural dispersion relation is present, i.e. when the ice is unbroken, while it disappears when the flexural component disappears, i.e. the floes get broken. This is the opposite of what could be expected from the floes array resonance mechanism discussed in zhao2017gap . Other possible mechanisms could involve, for example, edge waves, which can also participate in sustaining nonlinear resonant energy transfers marchenko1999parametric . Note that, to make the matter even more complex, nonlinear energy transfers through resonant triads could even happen across different kinds of waves phillips1981wave (edge waves being a possible candidate). There may also be other mechanisms that we have not yet imagined and that could also result in these high-frequency peaks.
@@
  • Ideally, in the field, one should perform detailed time-resolved and time-synchronized wave acceleration measurements in a wide area of sea ice, in order to be able to resolve the full dynamics of the two-dimensional ice sheet. This is a realistic attempt anno 2025, thanks to the low cost and accuracy of modern motion sensors such as the ISM330DHCX used in the OMB-v2021 (cost O(20 USD)), and the time accuracy provided by GPS Pulse Per Second (PPS) information (UTC second start can be determined with a jitter of at most O(10 ns) using GPS chips that cost O(30 USD). This was used, for example, in voermans2023estimating , which designed and deployed a low-cost open source logger that could reach microsecond accuracy in signal acquisition timestamps and logging rates of several kHz on a SD card, using a low-cost micro controller. With proper engineering, an open source SD-based wave in ice logger could be built with high-accuracy wave measurements and (sub-) microsecond timing accuracy for O(150 USD) including all material costs. This means that the hardware for a large-scale measurement campaign with 200 such loggers deployed as an array on the sea ice would cost O(30k USD), well within reach of a small- to medium-sized scientific project. This would enable a detailed analysis of the in-situ dispersion relation, spatial and spectral distributions of the energy, and of the effect of directionality.
  • In addition, detailed analytical calculations, as well as numerical simulations of the waves in ice equations using a 3-dimensional nonlinear solver and considering the effect of the full directional spectrum, could help to understand the energy transfers at stake constantin2018nonlinear .
  • Finally, it could be possible to investigate these phenomena in a laboratory experiment. However, this may be challenging due to (i) the need to properly scale the physical phenomena happening, which can result in several non-dimensional quantities conflicting with each others as it has previously been observed in the study of waves in ice attenuation rabault2019experiments , (ii) the likely need to consider fully 2-dimensional directional wave spectra, (iii) inherent challenges related to reflections and standing modes in a wave tank, especially when the effects to measure are subtle higher order effects that can easily be hidden by larger systematic errors induced by wave reflections, imperfect paddle-matching conditions, and finite-length effects of either the wavetank or the wave packets generated, as highlighter previously e.g. in rabault2016ptv ; marchenko2021laboratory .
@@
An interesting question is the impact and applicability of the present findings. While the present analysis gives us confidence that the spectral high-frequency peaks are real and may likely be related to interesting nonlinear dynamics, are these a physical and mathematical curiosity without much practical value, or are they important for sea ice dynamics or can be otherwise useful? If these high-frequency peaks could be reliably observed from satellite data, then they could possibly give information about the existence of flexural stresses in the ice and, hence, about the effective properties of the sea ice (in particular, through the information carried about the value of the $B$ factor in Eq. (2)). However, it remains to be seen whether these high-frequency peaks can be visible in satellite images and how much information can be inferred back.
Another interesting question is to determine how much additional energy dissipation can be induced by the nonlinear energy transfer to the high-frequency peak and the dissipation there. To estimate the importance of this effect, we can attempt to perform some order-of-magnitude calculations. To do so, we assume the following.
@@
Under these assumptions, the nonlinear energy transfer to the high-frequency peak and the dissipation there are in balance: $S_{\mathrm{nl}}=S_{\mathrm{diss},\mathrm{high-freq}}$. Therefore, estimating $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an estimate for $S_{\mathrm{nl}}$. As a consequence, estimating the relative importance between $S_{\mathrm{diss},\mathrm{main-peak}}$ (the energy dissipation at the main peak) and $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an indication of the fraction of the wave in ice attenuation in the main peak that comes from the nonlinear energy transfer and dissipation at higher frequency. To estimate this relative amount of energy dissipation between the main and the higher-frequency peaks, we use the following hypothesis.
@@
  • We assume that the wave energy spatial dissipation rate is similar to what is derived by yu2022 , in particular their Fig. 4.b which is based on a large set of test data: $k_{i}\propto f^{4.5}$.
@@
  • The temporal dissipation rate at a fixed location can be calculated as: $\beta=c_{g}k_{i}$ gaster1962note ; sutherland2017attenuation ; with deep water $c_{g}\propto f^{-1}$, this leads to $\beta\propto f^{3.5}$. Therefore, at a fixed location, the temporal dissipation rate ratio between the energy dissipation in the high-frequency peak and the main peak is $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=2^{3.5}\mathrm{PSD}_{\eta,2f}/\mathrm{PSD}_{\eta,f}$, where $\mathrm{DR}_{f}$ is the temporal dissipation rate at frequency $f$.
@@
  • $\mathrm{PSD}_{\eta}\propto f^{-4}\mathrm{PSD}_{\mathrm{acc}}$, and we are looking at $\mathrm{PSD}_{\mathrm{acc}}$ in our graphs, while the scaling formula for $k_{i}$ applies to $\mathrm{PSD}_{\eta}$; therefore, $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=(2^{3.5}/2^{4})\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc}f}\approx 0.7\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc},f}=0.7R_{f,2f}$, where $R_{f,2f}=\mathrm{PSD}_{acc,2f}/\mathrm{PSD}_{\mathrm{acc},f}$
@@
Taking into account the PSD plots presented in the figures above, $R_{f,2f}$ shows quite a large degree of variability. For some of the PSDs in the Antarctic $R_{f,2f}\approx O(1)$ ("A1"); for T15 $R_{f,2f}\approx O(0.1)$; for most other cases, we generally observe $R_{f,2f}\in{[O(0.1),O(0.01)]}$ ("O"). Therefore, our order-of-magnitude estimates based on the stationarity approximation indicate that damping in the high-frequency peak may represent up to 70% of the intrinsic energy dissipation for the main peak in the case "A1", 7% in the case T15, and 7% to 0.7% in the "O" cases. This indicates that this mechanism may, in certain circumstances, play a significant role (nearly doubling the amount of energy dissipation in the case "A1"), but that it may also in many cases be only a minor contribution to the total wave energy dissipation. However, note that these are only very coarse order-of-magnitude estimates. For example, we do not take into account the spectral width of the main vs. high-frequency peaks. Moreover, the high-frequency peak dissipation effect may already be somehow present in the wave in ice empirical dissipation laws that are fitted to observations, in which case the exponent in the frequency dependence of the energy dissipation may be biased. Therefore, additional and more detailed investigation is needed to provide a definitive conclusion, and one may need to consider a systematic, representative, large-scale, and well-sampled dataset to cover all typical cases of ice and wave conditions.

### v5

@@ 2.2.4 Bicoherence calculation
The (auto-, also known as self-) bicoherence (Mendel, 1991; He and Forbes, 2022) is used to estimate the degree of phase-locking between 3 frequencies $f_{1}$, $f_{2}$, and $f_{3}=f_{1}+f_{2}$ in a given timeseries signal. Here, we use the classical definition of the norm-2 bicoherence:
$b^{2}(f_{1},f_{2})=\frac{|\sum_{n}F_{n}(f_{1})F_{n}(f_{2})F_{n}^{*}(f_{1}+f_{2})|^{2}}{\sum_{n}|F_{n}(f_{1})F_{n}(f_{2})|^{2}\penalty 10000\ \sum_{n}|F_{n}(f_{1}+f_{2})|^{2}},$ (5)
@@ 3.1 Tempelfjorden 2015 (T15)
In order to further investigate whether the high-frequency spectral peak and bicoherence peak can correspond to a nonlinear wave interaction caused by wave triads, we compute the real-world wave dispersion relation by applying the cross-spectral density analysis between the VN1005 and VN1007 IMU sensors. These are the 2 IMUs that are located furthest away from each other, which gives a higher relative accuracy in the phase-shift measurements. Due to the separation distance, phase aliasing is observed. This phase aliasing is removed by enforcing a smooth monotonic increase in $\Delta\phi$ for increasing frequencies, by applying suitable corrections by $2\pi$ when aliasing jumps take place. The results are presented in Fig. 3, with the dispersion relation from Eqn. (3) included for different effective sea ice thickness values and otherwise standard landfast ice mechanical characteristics (we use similar values as Rabault et al. (2016b), i.e. $g=9.81$ m/s2, $E=2\times 10^{9}$ Pa, $\nu=0.3$, $\rho_{w}=1025$ kg/m3, $\rho_{i}=920$ kg/m3). As visible in Fig. 3, and in good agreement with the findings previously reported in Sutherland and Rabault (2016), a significant deviation from the open water dispersion relation, which matches well with a thin-plate elastic dispersion relation with the same typical thickness as measured in the field, is observed early in the data (up to around segment 4). We show the data points obtained about the low signal-to-noise ratio higher frequency components by using the low CSC threshold as thin crosses, and we expect that these are particularly noisy and should, therefore, only be taken as an indication. Later on, the dispersion relation collapses back to the deep-water dispersion relation, which we interpret to correspond to the development of cracks through the ice cover. Cracks were indeed observed when the instruments were recovered, while they were not present when the instruments were deployed, consistent with this interpretation.
@@
Figure 3: Empirical dispersion relation obtained from a cross-spectral analysis between the IMUs VN1005 and VN1007. Clear deviation from the deep water dispersion relation, in good agreement with the thin elastic plate dispersion relation and with previous results (Sutherland and Rabault, 2016), is observed until around segment 5. For later segments, the empirical dispersion relation matches the deep water one. Here, we show both the empirical dispersion relation obtained for the main PSD peak (large dots), and for the higher frequency components when these are present in the PSD (crosses; due to the low signal-to-noise ratio in the higher frequencies, these are obtained using a low value of the CSC threshold, and are therefore particularly noisy and should be taken only as an indication). The best-matching dispersion relation, which corresponds to an ice thickness of around 0.6 meters with the default ice parameters listed in the text (in good agreement with the field measurements reported in Sutherland and Rabault (2016)), provides the $B_{\mathrm{ref}}$ value for the $B$ coefficient in Eqn. (2).
@@
Knowing that triads can be expected, one can also observe, back on Fig. 1, that there are two significant dips in the main peak of the PSD of the IMUs furthest in (VN1002 and VN1007) relatively to the PSD of the IMU furthest out (VN1005). These dips are observed consistently over several FFT points, for both IMUs further in the ice, at frequencies around 0.150.15Hz and 0.190.19Hz, which sum up to the typical frequency of the high-frequency peak (around 0.340.34Hz). These may be a signature of the energy that is being "stolen" from the main peak and transferred into the high-frequency peak by the triad interactions.
Figure 4: Wave triad diagram corresponding to the thin plate dispersion relation that best matches the empirical dispersion relation observed for early segments in Fig. 3 (i.e. elastic plate with thickness 0.6m). This dispersion relation can sustain the wave triad conditions in the frequency range of the high-frequency peak that we observe. For the specific value of $\mathbb{k}_{3}$ corresponding to the PSD and bicoherence high-frequency peak, sets of locii for $\mathbb{k}_{1}$ and $\mathbb{k}_{2}$ that match the main peak in the PSD are identified (red). In this and similar figures, kxkx and $ky$ indicate the $x$ and $y$ components of a wave vector $\mathbb{k}=kx\mathbb{e}_{x}+ky\mathbb{e}_{y}$.
Since these data are obtained from field measurements, there are significant uncertainties on all parameters related to the ice properties (which are known to vary widely in time and space). In addition, the bicoherence peak is a relatively large "blob", and the PSD high-frequency peak is also relatively wide. Therefore, there is no single clear value for these parameters. Moreover, the "peak bicoherence triad" analyzed above corresponds only to the maximum of the bicoherence "blob". In reality, the spectra observed are broad-banded, and the high-frequency peak may rather be the consequence of a continuum of triad interactions than of a single triad. In particular, a full description of the dynamics would require integrating over all possible triads, akin to Hasselmann’s equation (Hasselmann, 1962). However, this is beyond the scope of the present work, and to offer a convincing comparison between these more advanced descriptions and observations, one may need significantly more detailed field data including measurements from a large array of instruments.
@@ 3.3 Spectra over Iridium
In addition to these observations, we also considered the PSD spectra from deployments in the MIZ in the case where small broken ice floes were encountered. These include in particular the outermost buoys in the dataset 2021-03 of Rabault et al. (2023), see more descriptions in Nilsen et al. (2021), and the outermost buoys in the datasets of Müller et al. (2025a, b). There, we could not find clear high-frequency peaks similar to those reported in Fig. 9. This further hints at the importance of having an ice sheet or a large enough ice floe to support significant flexural effects for observing the high-frequency peaks we report.
@@ 4 Discussion
In all cases, clear high-frequency peaks can be observed in some of the data segments. Since these high-frequency peaks are clearly above the noise background, are sometimes present and sometimes not (so they are not a systematic noise in the sensors and / or processing algorithms), and these can be observed over a range of instruments (VN100-based loggers, OMB-v2018 with a VN100 IMU, OMB-v2021 with an ISM330DHCX motion sensor, Sofar Spotter using GPS Doppler data) and deployments in various sea ice conditions and locations covering both the Arctic and the Antarctic, we can be confident that these are not noise nor a one-off coincidence, but a real recurring physical signal. Moreover, these peaks are relatively narrow isolated harmonics of the main energy peak and, therefore, most likely a different mechanism from the "rollover effect" that is sometimes reported in the literature (Wadhams et al., 1988) and may be associated with local wind energy injection or nonlinear effects (Li et al., 2017; Cheng et al., 2025), though debates have arisen in the community regarding the noise levels and validity of some of the observations (Thomson et al., 2021). Therefore, independently of the ongoing debate around the validity and presence of the rollover effect, we believe that our present observations of harmonic peaks in waves in ice energy data indicate a distinct signature and different mechanism.
@@
A critical question to elucidate regarding the high-frequency spectral peaks is whether these correspond to bound or free waves. Bound waves, which canonical example is the higher-order components of the Stokes solution (Lamb, 1924), are perturbations to the first-order wave solution that are present as the higher-order consequence of nonlinear terms. These can arise either from the effect of nonlinear terms on a single wave mode (as is the case with the higher order Stokes wave components), or as the effect of nonlinear terms on a superposition of wave modes (as the presence of nonlinear terms implies that the linear superposition of two or more wave modes results in an additional nonlinear residual term). Bound waves do not follow the dispersion relation, and are caused by and travel together with the main propagating wave, materializing as "disturbances" of the main wave. Free waves, on the contrary, propagate freely and follow the dispersion relation. Based on Fig. 3, it seems most likely that the high-frequency peak observed in the T15 dataset corresponds to a free wave that follows the dispersion relation. However, we acknowledge that the signal-to-noise ratio used there is low, so offering a definite confirmation that our high-frequency spectral peaks correspond to free waves would require additional observation and confirmation.
@@
While our observations seem to be compatible with the existence on nonlinear wave interaction and energy transfers between frequency modes through a wave triad mechanism, other mechanisms could also play a role and / or explain our observations. For example, Zhao et al. (2017) suggests that arrays of adjacent floating bodies (in our case, ice floes) can lead to nonlinear resonance and create high-frequency peaks in PSDs. However, this seems unlikely to be the cause of the observations in T15. Indeed, in this case, the high-frequency peak is observed when a flexural dispersion relation is present, i.e. when the ice is unbroken, while it disappears when the flexural component disappears, i.e. the floes get broken. This is the opposite of what could be expected from the floes array resonance mechanism discussed in Zhao et al. (2017). Other possible mechanisms could involve, for example, edge waves, which can also participate in sustaining nonlinear resonant energy transfers (Marchenko, 1999a). Note that, to make the matter even more complex, nonlinear energy transfers through resonant triads could even happen across different kinds of waves (Phillips, 1981) (edge waves being a possible candidate). There may also be other mechanisms that we have not yet imagined and that could also result in these high-frequency peaks.
@@
  • Ideally, in the field, one should perform detailed time-resolved and time-synchronized wave acceleration measurements in a wide area of sea ice, in order to be able to resolve the full dynamics of the two-dimensional ice sheet. This is a realistic attempt anno 2025, thanks to the low cost and accuracy of modern motion sensors such as the ISM330DHCX used in the OMB-v2021 (cost O(20 USD)), and the time accuracy provided by GPS Pulse Per Second (PPS) information (UTC second start can be determined with a jitter of at most O(10 ns) using GPS chips that cost O(30 USD)). This was used, for example, in Voermans et al. (2023), which designed and deployed a low-cost open source logger that could reach microsecond accuracy in signal acquisition timestamps and logging rates of several kHz on a SD card, using a low-cost micro controller. With proper engineering, an open source SD-based wave in ice logger could be built with high-accuracy wave measurements and (sub-) microsecond timing accuracy for O(150 USD) including all material costs. This means that the hardware for a large-scale measurement campaign with 200 such loggers deployed as an array on the sea ice would cost O(30k USD), well within reach of a small- to medium-sized scientific project. This would enable a detailed analysis of the in-situ dispersion relation, spatial and spectral distributions of the energy, and of the effect of directionality.
  • In addition, detailed analytical calculations, as well as numerical simulations of the waves in ice equations using a 3-dimensional nonlinear solver and considering the effect of the full directional spectrum, could help to understand the energy transfers at stake (Constantin, 2018).
  • Finally, it could be possible to investigate these phenomena in a laboratory experiment. However, this may be challenging due to (i) the need to properly scale the physical phenomena happening, which can result in several non-dimensional quantities conflicting with each others as it has previously been observed in the study of waves in ice attenuation (Rabault et al., 2019), (ii) the likely need to consider fully 2-dimensional directional wave spectra, (iii) inherent challenges related to reflections and standing modes in a wave tank, especially when the effects to measure are subtle higher order effects that can easily be hidden by larger systematic errors induced by wave reflections, imperfect paddle-matching conditions, and finite-length effects of either the wavetank or the wave packets generated, as highlighter previously e.g. in Rabault et al. (2016a); Marchenko et al. (2021).
@@
An interesting question is the impact and applicability of the present findings. While the present analysis gives us confidence that the spectral high-frequency peaks are real and may likely be related to interesting nonlinear dynamics, are these a physical and mathematical curiosity without much practical value, or are they important for sea ice dynamics or can be otherwise useful? Naturally, a first possible intrinsic impact is, as discussed above (and quantified below), the possible contribution of triads to indirectly increasing the attenuation rate of waves in ice. Another possible effect on the dynamics of the sea ice is that these may contribute to ice breakup, as discussed in Pierce et al. (2024). From a different viewpoint, considering a remote sensing perspective, if these high-frequency peaks could be reliably observed from satellite data, then they could possibly give information about the existence of flexural stresses in the ice and, hence, about the effective properties of the sea ice (in particular, through the information carried about the value of the $B$ factor in Eq. (2)). However, it remains to be seen whether these high-frequency peaks can be visible in satellite images and how much information can be inferred back.
We may attempt to determine how much additional energy dissipation can be induced by the nonlinear energy transfer to the high-frequency peak and the dissipation there. To do so, we can perform some order-of-magnitude calculations, assuming the following:
@@
Under these assumptions, the nonlinear energy transfer to the high-frequency peak and the dissipation there are in balance: $S_{\mathrm{nl}}=S_{\mathrm{diss},\mathrm{high-freq}}$. Therefore, estimating $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an estimate for $S_{\mathrm{nl}}$. As a consequence, estimating the relative importance between $S_{\mathrm{diss},\mathrm{main-peak}}$ (the energy dissipation at the main peak) and $S_{\mathrm{diss},\mathrm{high-freq}}$ provides an indication of the fraction of the wave in ice attenuation in the main peak that comes from the nonlinear energy transfer and dissipation at higher frequency. To estimate this relative amount of energy dissipation between the main and the higher-frequency peaks, we use the following hypothesis:
@@
  • We assume that the wave energy spatial dissipation rate is similar to what is derived by Yu et al. (2022), in particular their Fig. 4.b which is based on a large set of test data: $k_{i}\propto f^{4.5}$.
@@
  • The temporal dissipation rate at a fixed location can be calculated as: $\beta=c_{g}k_{i}$ (Gaster, 1962; Sutherland et al., 2017a); with deep water $c_{g}\propto f^{-1}$, this leads to $\beta\propto f^{3.5}$. Therefore, at a fixed location, the temporal dissipation rate ratio between the energy dissipation in the high-frequency peak and the main peak is $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=2^{3.5}\mathrm{PSD}_{\eta,2f}/\mathrm{PSD}_{\eta,f}$, where $\mathrm{DR}_{f}$ is the temporal dissipation rate at frequency $f$.
@@
  • $\mathrm{PSD}_{\eta}\propto f^{-4}\mathrm{PSD}_{\mathrm{acc}}$, and we are looking at $\mathrm{PSD}_{\mathrm{acc}}$ in our graphs, while the scaling formula for $k_{i}$ applies to $\mathrm{PSD}_{\eta}$; therefore, $\mathrm{DR}_{2f}/\mathrm{DR}_{f}=(2^{3.5}/2^{4})\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc}f}\approx 0.7\mathrm{PSD}_{\mathrm{acc},2f}/\mathrm{PSD}_{\mathrm{acc},f}=0.7R_{f,2f}$, where
    $R_{f,2f}=\mathrm{PSD}_{acc,2f}/\mathrm{PSD}_{\mathrm{acc},f}$ .
@@
Taking into account the PSD plots presented in the figures above, $R_{f,2f}$ shows quite a large degree of variability. For some of the PSDs in the Antarctic $R_{f,2f}\approx O(1)$ (case "A1"); for T15 $R_{f,2f}\approx O(0.1)$ (case T15); for most other cases, we generally observe $R_{f,2f}\in{[O(0.1),O(0.01)]}$ (case "O"). Therefore, our order-of-magnitude estimates based on the stationarity approximation indicate that damping in the high-frequency peak may represent up to 70% of the intrinsic energy dissipation for the main peak in the case "A1", 7% in the case T15, and 7% to 0.7% in the "O" cases. This indicates that this mechanism may, in certain circumstances, play a significant role (nearly doubling the amount of energy dissipation in the case "A1"), but that it may also in many cases be only a minor contribution to the total wave energy dissipation. However, note that these are only very coarse order-of-magnitude estimates. For example, we do not take into account the spectral width of the main vs. high-frequency peaks. Moreover, the high-frequency peak dissipation effect may already be somehow present in the wave in ice empirical dissipation laws that are fitted to observations, in which case the exponent in the frequency dependence of the energy dissipation may be biased. Therefore, additional and more detailed investigation is needed to provide a definitive conclusion, and one may need to consider a systematic, representative, large-scale, and well-sampled dataset to cover all typical cases of ice and wave conditions.